y=-2/5 x-18
2x-3y=22

Answer by checkley77(12569)

y=-2/5x-18
2x-3y=22
2x-3(-2x/5-18)=22
2x+6x/5+54=22
(10x+6x)/5=22-54
16x/5=-32
16x=5*-32
16x=-160
x=-160/16
x=-10 answer.
y=-2/5*-10-18
y=20/5-18
y=4-18
y=-14 answer.
proof:
2*-10-3*-14=22
-20+42=22
22=22